OpenStudy (anonymous):

The length of a diagonal of a square is 7 inches. Find the perimeter of the square. A. 25 inches B. 12 inches C. 19.6 inches D. 28 inches

OpenStudy (goldphenoix):



OpenStudy (goldphenoix):

Do you know how to set up the equation?

OpenStudy (anonymous):

2a^2=7^2

OpenStudy (goldphenoix):

Not quite. You're missing b. So the equation would look like: \[\large \large a^2 + b^2 = 7^2\]

OpenStudy (goldphenoix):

Now simplify. What would 7^2 become?

OpenStudy (anonymous):

49

OpenStudy (goldphenoix):

So it would look like: \[\large \large a^2 + b^2 = 49\]

OpenStudy (goldphenoix):

Since its a square, you can write it as: \[\large \large a^2 + a^2 = 49\] Now what does a^2 + a^2 give us?

OpenStudy (anonymous):

2a^2.

OpenStudy (anonymous):

@RPguy @eSpeX

OpenStudy (espex):

You have worked your equation down to \(2a^2=49\), isolate your \(a^2\) next.

OpenStudy (espex):

How would you get your \(a^2\) alone?

OpenStudy (anonymous):

\[\sqrt{a}^{2}=\sqrt{49}\]

OpenStudy (espex):

Close, but recall that you have \(2a^2\), so you will need to divide out the 2 before you can take the square root and get 'a'.

OpenStudy (anonymous):

\[\sqrt{24.5}=\sqrt{a}^{2}\]

OpenStudy (espex):

Yes, making 'a' equal to ?

OpenStudy (espex):

And then you know that the perimeter of a square is equal to \(4a\)

OpenStudy (anonymous):

so it will be c

OpenStudy (espex):

Yes, nice work. :)

OpenStudy (anonymous):

thank you :)

OpenStudy (espex):

You're welcome.